OpenStudy (anonymous):

Find the energy of a photon of light if the wavelength of the light wave is 6.88 x 10^–7 m. A. 2.89 x 10^–19 J B. 4.56 x 10^–40 J C. 3.46 x 10^18 J D. 7.56 x 10^–20 J

OpenStudy (chmvijay):

use E= hv =hc/lambda lambda is wavelength h- planks constant c velocity of light in vacuum

OpenStudy (anonymous):

Here Lambda λ= 6.88*10^-7 m E=hv or E=hc/ λ E=(6.63*10^-34)(3*10^8)/6.88*10^-7) E=2.89*10^-19 J ( Option A) What if I tell you a trick to solve it Fast just Solve the Exponents(Powers): E= 10^-34 * 10^8 /10^-7 or E= 10^-34+8+7 E=10^-19 and Multiply the Coefficients 6.63*3/6.88 = 2.89